# Stability of Continuous Time Quantum Walks in Complex Networks

### v2

@@ IV.5 Quantum-Classical Distance
The quantum-classical distance is a measure that tells the divergence between quantum and classical evolution. If it is 0, then the quantum evolution is fully classicalized, which is not desired, and if $1$, where the system evolves entirely different from classical evolution, indicates that decoherence has not classicalized the system and quantum properties are preserved. Quantum classical distance for complete cycle and star topologies is already studied in [bressanini2022decoherence].
Using these metrics on different topologies, we analyze which topology is more preferred under different decoherence.
V Results
V.1 CTQW on Complete, Cycle and Star Networks
Figure 2: Effect of network size and decoherence rate on the $\ell_{1}$-norm of coherence for a star graph with central-hub initialization. Panels (a)–(c) show the coherence decay for different network sizes NN under (a) Haken–Strobl noise, (b) QSW dynamics, and (c) intrinsic decoherence, each evaluated at fixed decoherence rates ($\gamma=0.1$ or $p=0.1$). (d)–(f) The bottom row illustrates the coherence decay for increasing decoherence rates ($\gamma$ or $p$) for the three corresponding decoherence models on a network of fixed size $N=10$.
@@ V.1 CTQW on Complete, Cycle and Star Networks
Figure 3: Probability distribution of the CTQW on some simple network topologies with network size N = 10. To examine the probability distribution of CTQW on various network structures, we consider three representative topologies: cycle, star, and complete graphs, each consisting of $N=10$ nodes. In each network, the red-colored node indicates the site where the walker is initially localized with probability $1$ at time $t=0$. The temporal evolution of the probability distribution is illustrated under four scenarios: (a)–(d) the noiseless case, (e)–(h) Haken–Strobl noise, (i)–(l) intrinsic decoherence, and (m)–(p) QSW. The four columns correspond to the different network configurations: cycle (first column), star with hub node initialization (second column), star with peripheral node initialization (third column), and complete graph (fourth column).
To study the evolution of the CTQW on various network topologies, we first examine the dynamics on simple networks such as the cycle, complete, and star topologies. Star networks consists of two types of nodes: peripheral nodes, each with degree $1$, and a central hub node with degree N1N-1, so for this network, we consider two distinct initial conditions: in the first case, the walker is localized at the central hub node, and in the second case, it is initialized on one of the peripheral nodes. In contrast, both cycle and complete networks consist of nodes of uniform degree. Therefore, for these topologies, the CTQW is initially located on a randomly selected node due to the symmetry of the graph. In all the cases, the initially localized node refers to the node with probability 1 at the initial time step.
To probe the system’s stability, we first conduct a detailed scaling analysis of the star network with central node initialization, studying how its $\ell_{1}$-norm of coherence depends on both network size N and decoherence rate. As shown in Fig. 2(a), for Haken-Strobl noise the time taken for coherence decay increases with network size, showing that stability improves as $N$ increases. For QSW, however, the opposite is true. Fig. 2(b) shows that as $N$ increases, the coherence decays faster, indicating that larger networks are little less stable under this decoherence. For intrinsic decoherence, Fig. 2(c) shows that as $N$ increases, the coherence saturates faster and to a higher steady-state coherence value, suggesting greater stability. Although smaller networks exhibit a slight peak before stabilizing, the reduced overall decay for larger $N$ shows that stability is enhanced. Next, we study the dependence on decoherence rates for a fixed network size of $N=10$. As shown in Fig. 2(d) for Haken-Strobl noise and Fig. 2(e) for QSW, stability uniformly decreases as the respective rates $\gamma$ and $p$ increase. For intrinsic decoherence Fig. 2(f) shows that while larger $\gamma$ value dampen initial oscillations and lead to faster saturation, the final steady-state coherence value is dependent on the network size, not on the decoherence rate. Having studied this baseline scaling, we now proceed to a detailed comparison of the dynamics across the complete, cycle, star topologies with fixed size $N=10$. For this and all subsequent simulations, fixed decoherence rates of $\gamma=0.1$ and $p=0.1$, as this scaling analysis confirms they are large enough to make the decoherence effects clearly visible.
In the absence of noise, the CTQW on the cycle network as shown in Fig. 3(a), explores not only the initially localized node but also other nodes in the network. In contrast, for both the star and complete network topologies, the CTQW tends to remain localized near the initially occupied node, maintaining a high occupation probability at that node, as illustrated in Fig. 3(b),(d). As shown by Razzoli et al. [razzoli2022universality], the dynamics of a CTQW initialized on a fully connected hub (a node with degree N-1) are universal, meaning they are determined only by the size of the network, N, and are independent of the rest of the graph’s structure. Our simulations shows that this universality is not only true for noisless case, but also true for Haken-Strobl and intrinsic decoherence as shown in Fig. 3(f),(h),(j),(l). Moreover, the probability oscillations in the complete graph and the star network with hub node initialization are identical and exhibit a higher frequency compared to the star network with peripheral node initialization. For the star network with initial localization at a peripheral node in Fig. 4(c), the $\ell_{1}$-norm of coherence oscillates with lower frequency compared to the cases of the star with central node initialization and the complete graph, both of which show similar and more rapid oscillatory behavior as shown in Fig. 4(b),(d). This observation is consistent with previous findings by Xu et al. [xu2009exact], who showed that transition probabilities are high-frequency revivals in the case of hub node excitation, along with strong long-term localization at the initial site for the star network. In addition to it, that study also demonstrated that the transport dynamics on a star graph with hub node excitation are equivalent to those on a complete graph of the same size. In contrast, for the cycle network, the $\ell_{1}$-norm of coherence displays fluctuations that do not go to zero, as shown in Fig. 4(a) and it exhibits a higher coherence, as measured by the $\ell_{1}$-norm of coherence, than both the complete and star networks. This can be attributed to the absence of high-degree nodes in the cycle network, which allows for delocalized dynamics and a higher value of coherence. These results suggest that the dynamics of CTQW are highly sensitive to both the underlying network structure and the choice of initial localization. Under Haken–Strobl noise, the occupation probabilities of the CTQW on the cycle graph as shown in Fig. 3(e) converge to a uniform distribution significantly faster than on the other topologies given in Fig. 3(f)-(h), indicating rapid delocalization and the onset of a steady state. This behavior resembles the equilibration of probabilities observed in classical continuous-time random walks [mulken2011continuous]. The system’s convergence to the maximally mixed state $\frac{\mathbb{I}}{N}$ (equilibration of probabilities and $\ell_{1}$-norm of coherence decays to zero) is consistent with the fact that both the Haken-Strobl noise and QSW Liouvillians are unital and Davies irreducible [davies1976quantum] for connected graphs. As detailed in Appendix  A.1 and  A.2, we verify irreducibility using the algebraic criteria of Zhang and Barthel [zhang2024criteria], which guarantees a unique and maximally mixed steady state. The quantum coherence, quantified by the $\ell_{1}$-norm of coherence, also decays much more rapidly on the cycle topology as shown in Fig. 4(a) compared to the star and complete networks in Fig. 4(b)-(d). To determine the rate of coherence loss, we fitted the simulated coherence data using the stretched exponential Kohlrausch function [kohlrausch1854theorie]:
@@
Here, $\lambda$ is the effective decay rate that determines how quickly the function decreases (increases) over time, while $\beta$ is the stretching exponent that characterizes the nature of the decay. A value of $\beta=1$ corresponds to standard exponential decay, whereas $\beta<1$ coherence decay slower than exponential decay and for $\beta>1$ the quantum walker loses coherence faster than exponential decay. However, the overall behavior of the coherence decay is determined by the combined influence of both $\lambda$ and $\beta$, which can be calculated as a single characteristic relaxation time, $\tau$, which represents the mean lifetime of the coherence from its normalized initial value. given by  [gueguen2015relationship]:
@@
We fitted coherence data obtained for the Haken-Strobl noise and QSW as shown in Fig. 4(e),(f),(g),(h) and the extracted coefficients $\lambda$, $\beta$ and their characteristic relaxation time, summarized in Table 1. These characteristic relaxation time $\tau$, further support our observations: the star network with the central node as the initial localization and the complete network exhibit slower coherence decay characterized by higher and identical $\tau$, indicating greater robustness against Haken–Strobl decoherence compared to both the cycle network and the star network initialized at a peripheral node. Moreover this identical $\lambda$ and $\beta$ as shown in Table 1 for the fully connected, initially localized node, show that the universality discussed earlier holds for Haken-strobl noise, applying not only to probability but to coherence as well. Among all the network configurations considered, the cycle network performs the worst in maintaining quantum coherence for a long time and the star network with peripheral node initialization in a short time. However, it should be noted that for Haken-Strobl decoherence, all network topologies exhibit $\beta<1$ as in Table 1, indicating that the decay of coherence is slower than exponential in all cases.
@@
The evolution of CTQW under intrinsic decoherence is shown in Fig. 3(i),(j),(k),(l), respectively, on cycle, star, and complete networks. From these figures, we observe distinct differences in the localization behavior, for instance, on star and complete topologies, CTQW tends to remain localized with a significantly high probability at the initial node compare than at other nodes in the network. In contrast, for the cycle topology, although the probabilities do not become exactly uniform, there is no strong localization on the initially localized node, unlike in star and complete topologies. Instead, the occupation probabilities of different nodes exhibit persistent oscillations over time (see Fig. 3(i)). As shown in Fig. 4(a)-(d) across all network types, coherence under intrinsic decoherence is preserved to varying extents. In the cycle network, the $\ell_{1}$-norm of coherence exhibits temporal fluctuations even at large times. For the star network with initial localization at a peripheral node, coherence exhibits damped oscillations before settling into a steady state value. Interestingly, for both the complete network and the star network with central hub node initialization, the coherence also converges to the same steady state value observed in the peripheral initialization case. Although the cycle network retains some coherence under intrinsic decoherence, the level of preservation is lower compared to the star and complete topologies. Notably, the cycle exhibits oscillatory behavior in both probability and coherence dynamics under intrinsic decoherence, a feature that is absent in the star and complete networks.
In the case of QSW decoherence, the decay of coherence is faster for complete network than other topologies. However, decay of coherence under QSW decoherence is much faster than Haken-Strobl for all network topologies, as shown in Table 1 and the stability ranking between topologies is also inverted. Star graph with peripheral node initialization takes longer to reach the same probability distribution and exhibits a slower decay of coherence compared to other topologies and the star graph with center initialization, as seen in Fig. 3(m)-(p) and Fig. 4(e)-(h). This observation is further supported by $\tau$ as shown in Table 1, which shows that the coherence takes a longer time to decay in the star topology with initialization on peripheral node than all the other topologies, indicating greater stability of this topology under QSW induced decoherence. This complete reversal suggests that stability under QSW is inversely related to node connectivity. Because QSW is a link-based decoherence model (Eq. (4)), it appears to more dissipation as the connections in the network increases, explaining why the highly-connected complete graph and star-hub configurations are so fragile, while the minimally connected peripheral node is the most stable.
@@
The behavior of quantum-classical distance, fidelity with the initial state, and von Neumann entropy is presented in Appendix B for networks of size $N=10$ Fig. 9 and Fig. 10. Under Haken-Strobl noise and intrinsic decoherence, the star network (with hub node initialization) and the complete graph exhibit high stability, indicated by elevated quantum-classical distance and fidelity, along with low entropy for Haken-Strobl. For the QSW, the star network initialized at a peripheral node shows improved stability, as seen from its high fidelity and low entropy. Interestingly, the quantum-classical distance highlights the cycle network as especially robust against classicalization.
V.2 CTQW on Complex Network Topologies
@@
We extend our analysis to more complex topologies to investigate how different network structures influence the stability of quantum systems. Specifically, we study the evolution of CTQW on Erdős–Rényi, small-world, and scale-free networks. As the degree of the initially localized node can significantly affect the dynamics (see Appendix C Fig. 11), we choose a high-degree node as the initial localization in the heterogeneous scale-free network. In contrast, the Erdős–Rényi and small-world networks have nodes with approximately similar average degrees, so the impact of the initial node selection is expected to be comparatively minimal in these cases. In all cases, the initially localized node is defined as the node with probability $1$ at the initial time step (see Fig. 5).
Erdős–Rényi Network
Refer to captionProbability
Small-World Network
Refer to captionProbability
Scale-Free Network
Refer to captionProbability
@@
Fig. 5 shows that the scale-free network exhibits distinct behavior compared to the Erdős–Rényi and small-world networks, particularly in the noiseless case, as illustrated in Fig. 5(a),(e),(i). Node centralities in scale-free networks vary significantly due to their inherently heterogeneous structure. As a result, the evolution of CTQW on such networks depends strongly on the centrality of the node where the walker is initially localized. For instance, when the initially localized node has the highest degree but not the highest closeness centrality, the occupation probabilities at the highest-degree and highest-closeness nodes exhibit an anti-phase synchronization pattern (see Appendix C Fig. 11). The probability amplitudes of anti-phase synchronized nodes vary depending on the specific configuration of the network structure. Notably, strong localization at the initially occupied hub node is observed in heterogeneous networks, such as, the scale-free network Fig. 5(i) and star networks. This behavior is consistent with the previous study by Xu et al. [xu2009exact] who showed that, in the thermodynamic limit ($\lim{N\to\infty}$), transition probabilities in star networks tend to remain localized at the initially localized site. In contrast, Erdős–Rényi and small-world networks lack pronounced degree heterogeneity and central hubs, leading to a more delocalized evolution of the CTQW. Interestingly, however, Fig. 6 shows that, in the absence of decoherence, both small-world and Erdős–Rényi networks maintain a higher $\ell_{1}$-norm of coherence compared to the scale-free network. This highlights the influence of network structure, suggesting that while hubs facilitate localization, they may simultaneously contribute to reduced coherence in noiseless condition.
@@
Under the influence of Haken–Strobl noise, the occupation probability at the initially localized node decays rapidly in both the Erdős–Rényi and small-world networks Fig. 5(b),(f), eventually leading to a uniform distribution. This renders the initially localized node indistinguishable from others, indicating complete delocalization across the network. In contrast, the scale-free network does not exhibit complete delocalization in short time Fig. 5(j). Although the occupation probability at the initially localized node decays over time, the distribution remains uneven and converges toward uniformity at a significantly slower rate than that observed in the Erdős–Rényi and small-world networks.
Furthermore, we extracted the coherence decay coefficients $\lambda$ and $\beta$ by fitting the coherence data as shown in the red curves in Fig. 6 using Eq. (11). From these, we then calculated the relaxation time, $\tau$ using Eq. 12, and all three parameters ($\lambda$$\beta$ and $\tau$) are summarized in Table 2. The red curves in Fig. 6(c) show that the scale-free network preserves quantum coherence for a longer duration under Haken–Strobl noise, as indicated by its relatively slow decay, showed by a high relaxation time. In the presence of intrinsic decoherence, the CTQW exhibits partial localization at the initially localized node in the scale-free network Fig. 5(k), similar to the behavior observed in the noiseless case, but with a fixed probability. Localization behavior of CTQW is not observed in the Erdős–Rényi and small-world networks, as illustrated in Fig. 5(c),(g). Interestingly, the $\ell_{1}$-norm of coherence is preserved under intrinsic decoherence for any of the network topologies. This suggests that, despite inducing partial localization in the scale-free network, intrinsic decoherence preserves quantum coherence across all topologies, ensuring stability for all topologies. On the other hand, QSW decoherence lead to a complete delocalization of the quantum walker across all network types. As shown in Fig. 5(d),(h),(l), node probabilities tend to equalize, erasing the initial localization, particularly evident in the scale-free network, where the initial localization is entirely lost. Relaxation time ($\tau$) from Table 2, and purple curves in Fig. 6 further shows that QSW induces the fastest coherence decay among all the decoherence. Although Erdős–Rényi networks retain coherence for more time compared to other topologies under QSW, overall coherence degrades rapidly across all complex network topologies. However, this decay is not faster than exponential as in simple topologies discussed in Sec. V.1, as indicated by $\beta<1$.
Decoherence Network $\lambda$ $\beta$ $\tau$
Haken-Strobl Erdős-Rényi 0.10 0.84 10.96
Small-World 0.17 0.96 5.99
Scale-Free 0.24 0.37 17.43
Quantum Stochastic Erdős-Rényi 0.71 0.45 3.49
Small-World 1.53 0.90 0.69
Scale-Free 0.96 0.62 1.50
@@
Table 2: Coherence decay fitting parameters for Erdős-Rényi, small-world and scale-free under Haken-Strobl and QSW decoherence.
@@
Further to assess the system’s stability, we utilize a set of standard metrics, quantum-classical distance, von Neumann entropy, and fidelity. The quantum-classical distance in Fig. 9(a)-(d) tells that the scale-free network exhibits greater resistance to classicalization under decoherence compared to the Erdős–Rényi and small world network, since as time progresses, scale-free has higher quantum classical distance than other topologies underlying all decoherence dynamics. It shows a slower decay of quantum-classical distance in the presence of both QSW and Haken-strobl both of which classicalizes the system over long timescales. For intrinsic decoherence in Fig. 9(c), the distance does not decay to zero, indicating that classicalization does not occur and quantum features are consistently preserved for all networks. For the noiseless scenario, across all network topologies, the quantum-classical distance converges to the theoretical value $\lim_{t\to\infty}D_{QC}(t)=1-\frac{1}{N}$ [gualtieri2020quantum], as represented by the dashed line in Fig. 9(a).
@@
The fidelity plots in Fig. 9(e)-(h) show that the scale-free network topology exhibits localization near the initial state since it has high fidelity with the initial state. Higher the fidelity lesser the evolved state has deviated from the initial state. On contrast, the Erdős-Rényi and small-world networks have a little less fidelity across various decoherence.
@@
Von Neumann entropy plots Fig. 10(a)-(c) show that entropy is high for Erdős-Rényi in all decoherence. Entropy is least for scale-free in the case of Haken-Strobl and QSW, showing scale-free is not much deviated from the pure state as compared to Erdős-Rényi. This implies that scale-free is more resilient to all noises and Erdős-Rényi is more prone to mixed states than other topologies. For the noiseless case, the entropy always remains zero.
VI Conclusion

### v3

@@ II.3.5 Quantum-Classical Distance
The quantum-classical distance is a measure that tells the divergence between quantum and classical evolution. If it is 0, then the quantum evolution is fully classicalized, which is not desired, and if $1$, where the system evolves entirely different from classical evolution, indicates that decoherence has not classicalized the system and quantum properties are preserved. Quantum classical distance for complete, cycle and star topologies is already studied in [16].
Using these metrics on different topologies, we analyze which topology is more preferred under different decoherences.
III Results
III.1 CTQW on Complete, Cycle and Star Networks
@@
Figure 2: Probability distribution of the CTQW on simple network topologies. To examine the probability distribution of CTQW on various network structures, we consider three representative topologies: cycle, star, and complete graphs, each consisting of $N=10$ nodes. In each network, the red-colored node indicates the site where the walker is initially localized with probability 11. The temporal evolution of the probability distribution is illustrated under four scenarios: (a)–(d) the noiseless case, (e)–(h) Haken–Strobl noise, (i)–(l) intrinsic decoherence, and (m)–(p) QSW. The columns from left to right correspond to the cycle, star (hub initialization), star (peripheral initialization) and complete network.
To investigate the CTQW dynamics on different network topologies, we first analyze the probability distribution of the CTQW on each node for the cycle, complete, and star networks. For the star network, we consider two cases; in one case, the walker is initialized at the hub, and in another case, it is initialized at any peripheral node. For the cycle and complete networks, owing to degree symmetry, it starts from a randomly selected node. In all cases, the initially localized node has probability $1$ at $t=0$. To assess system stability, we additionally compute the $\ell_{1}$-norm of coherence, the quantum–classical distance, the fidelity with the initial state, and the von Neumann entropy. For all analyzes, we fix the network size at $N=10$, as the $\ell_{1}$-norm of coherence for star networks of different sizes (with hub initialization) exhibits qualitatively similar behavior; second, large-sized networks preserve coherence longer under Haken-Strobl and intrinsic decoherence (see Fig. 8(a)-(c) in Appendix A). The decoherence parameters are set to $\gamma=0.1$ and $p=0.1$, since these values are sufficiently large to make decoherence effects clearly visible (Fig. 8(d)-(f)).
Temporal probability plots (Fig. 2) show that, in the absence of noise, the CTQW on the cycle network (Fig. 2(a)) explores not only the initially localized node but also other nodes in the network. In contrast, for both the star and complete networks (Fig. 2(b)-(d)), the walker remains strongly localized at the initial node, maintaining a high occupation probability. The identical behavior of the hub-initialized star and complete networks, owing to the fully connected (N1N-1) initial node is consistent with the universality reported by Razzoli et al. [48], who showed that, in the noiseless case, CTQW dynamics initiated at a fully connected hub depend only on the network size and not on the remaining graph structure. Notably, our results indicate that this universality is not only true for noiseless case, but true for Haken-Strobl and intrinsic decoherence as shown in Fig. 2(f),(h),(j),(l).
We next analyze the temporal behavior of the $\ell_{1}$-norm of coherence (Fig. 3). We observe that the cycle network exhibits persistent oscillations that do not decay to zero (Fig. 3(a)) and has the highest coherence among all topologies. This reflect that network topologies that don’t have hub nodes promotes delocalized dynamics and enhanced coherence, thereby reflecting a trade-off between localization and coherence in the noiseless case. In contrast, the complete network and the hub-initialized star network (Fig. 3(b),(d)) display higher-frequency oscillations in both occupation probabilities and coherence than the peripheral-initialized star network (Fig. 3(c)), consistent with the high-frequency revivals in transition probabilities reported by Xu et al. [49]. These results highlight the strong dependence of CTQW dynamics on network structure and initial conditions.
Under Haken-Strobl noise (Fig. 2(e)-(h)), all networks converge to a uniform probability distribution, with the cycle network reaching the steady state fastest. This behavior resembles the equilibration of probabilities observed in classical continuous-time random walks [27]. The simultaneous decay of coherence to zero (Fig. 3) and the equilibration of probabilities shows the onset of maximally mixed state $\frac{\mathbb{I}}{N}$ and is consistent with the fact that both the Haken-Strobl noise and QSW Liouvillians are unital and Davies irreducible [50] for connected graphs. As shown in Appendix B.1 and Appendix B.2, irreducibility of the Liouvillians is verified using the algebraic criteria of Zhang and Barthel [51], ensuring a unique maximally mixed steady state.
Under intrinsic decoherence (Fig. 2(i)-(l)), the star and complete networks remain strongly localized at the initial node and rapidly approach a steady state. In contrast, the cycle topology does not exhibit strong localization; instead, node occupation probabilities display damped oscillations, and the $\ell_{1}$-norm of coherence shows persistent temporal fluctuations even at long times (Figs. 2(i) and 3(a)). Although coherence survives across all topologies (Fig. 3(a)-(d)), it is highest for the star and complete networks than cycle network. For peripheral initialization in the star network (Figs. 2(k) and 3(c)), coherence and initially localized node probability exhibits damped oscillations and a slower relaxation compared to hub initialization, while ultimately converging to the same steady-state value as for complete network.
Under Haken-Strobl and QSW decoherence, the coherence decays with time. To quantify the decay rate of the $\ell_{1}$-norm of coherence, we fit the simulated data using the stretched-exponential Kohlrausch function [52]:
@@
Here, $\lambda$ is the effective decay rate that determines how quickly the function decreases over time, while $\beta$ is the stretching exponent that characterizes the nature of the decay. However, the overall behavior of the coherence decay is determined by the combined influence of both $\lambda$ and $\beta$. This can be quantified as a single characteristic relaxation time, $\tau$, which represents the mean lifetime of the coherence from its normalized initial value [53]:
@@
We fitted coherence data obtained for the Haken-Strobl noise and QSW as shown in Fig. 3(e)-(h) and the extracted coefficients $\lambda$, $\beta$ and $\tau$ summarized in Table 1. The complete and hub initialized star networks exhibit slower coherence decay characterized by the highest $\tau$, followed by the star peripheral initialized case, with the cycle network showing the fastest decay in the case of Haken-Strobl noise. We also observe that coherence decays significantly faster under QSW decoherence than under Haken–Strobl for all network topologies, and the stability ranking correspondingly inverted. Specifically, under QSW, the star graph with peripheral node initialization take longer to equalize probability than the star with hub initialization and the complete network, whereas the opposite trend is observed under Haken–Strobl decoherence. Notably, the peripheral node initialized star and cycle network exhibits comparatively slower coherence decay under QSW, indicating enhanced robustness. Since QSW is a link-based decoherence model (Eq. (4)), dissipation increases with the number of connections, rendering highly connected structures such as the complete graph and star-center configurations more fragile, while minimally connected configurations, such as peripheral-node initialization, remain more stable. To further elucidate the origin of these stability hierarchies, Appendix C establishes a direct correlation between the relaxation time $\tau$ and specific properties of the initialized node. These results suggest that the closeness centrality is important to capture actual $\tau$ for the simple networks under Haken-Strobl decoherence and inverse density weighted degree for the QSW decoherence. Overall our result suggest that in single CTQW, the preservation of coherence and stability depends on the network structure, choice of initial localization and decoherence types.
@@
Further, the stability of CTQW is systematically analyzed in Appendix D through quantum-classical distance, fidelity with the initial state (Fig. 10), and von Neumann entropy (Fig. 11). Our result show that under Haken-Strobl noise and intrinsic decoherence, the star network (with hub node initialization) and the complete graph exhibit high stability, indicated by elevated quantum-classical distance and fidelity, along with low entropy for Haken-Strobl noise. For the QSW, the star network initialized at a peripheral node shows improved stability, as seen from its high fidelity, low entropy and higher value of quantum-classical distance.
In summary, star network with peripheral initialization shows greater stability under QSW decoherence. However, for all other decoherence models, the complete graph and the star network with central node initialization outperform other topologies offering enhanced stability. Furthermore, the cycle network exhibits resistance to intrinsic decoherence induced steady state even at long times.
III.2 CTQW on Complex Network Topologies
@@
We extend our analysis to more complex topologies to investigate how different complex network structures influence the stability of quantum systems. Specifically, we study the evolution of CTQW on Erdős–Rényi, small-world, and scale-free networks, maintaining a fixed average degree of $\langle k\rangle=4$ and network size $N=100$; the broader impact of varying average connectivity and network size is detailed in Appendix E (Fig. 12 and Fig. 13). As the degree of the initially localized node can also significantly affect the dynamics (see Appendix F Fig. 14), we choose a high-degree node as the initial localization in the heterogeneous scale-free network. In contrast, the Erdős–Rényi and small-world networks have nodes with approximately similar average degrees, so the impact of the initial node selection is expected to be comparatively minimal.
Erdős–RényiSmall–WorldScale–FreeRefer to captionProbabilityRefer to captionProbabilityRefer to captionProbability
@@
Fig. 4 shows strong localization at the initially occupied hub node in the scale-free network. In contrast, Erdős–Rényi and small-world networks lack pronounced degree heterogeneity and central hubs, leading to a more delocalized evolution of the CTQW. Scale-free networks are heterogeneous, so the evolution of CTQW also strongly depends on the centrality of the node where the walker is initially localized (see Appendix F Fig. 14). Depending on the network configuration, a given node may or may not simultaneously possess multiple centrality maxima. For example, a node may have both the highest degree and the highest closeness centrality. However, when these maxima occur at different nodes and the walker is initialized at one of them, the occupation probabilities of walker at the highest-degree and highest-closeness nodes exhibit an anti-phase synchronization pattern. When the initialized node has both the highest degree and closeness centrality, the dynamics exhibit strong localization at that node, except in the presence of QSW decoherence. Xu et al. [49] demonstrated that, in the thermodynamic limit ($N\to\infty$), transition probabilities in star networks remain localized at the initial site in noiseless case. Since both star and scale-free networks are heterogeneous and contain hub nodes, our results suggest that localization at a centrally initialized node is a general feature of such networks, valid not only in the noiseless case but also under intrinsic decoherence. Further, Fig. 5 shows that, in the absence of decoherence, both small-world and Erdős–Rényi networks maintain a higher $\ell_{1}$-norm of coherence compared to the scale-free network. This highlights the influence of network structure, suggesting that while hubs facilitate localization, they may simultaneously contribute to reduced coherence in noiseless conditions.
@@
Under the influence of Haken–Strobl noise, the occupation probability at the initially localized node decays rapidly in both the Erdős–Rényi and small-world networks Fig. 4(b),(f), eventually leading to a uniform distribution. This renders the initially localized node indistinguishable from others, indicating complete delocalization across the network. This process is accompanied by the $\ell_{1}$-norm of coherence decaying to zero (Fig. 5(a),(b)), making the transition to a steady state (Appendix B.1). In contrast, the scale-free network does not exhibit complete delocalization in short time Fig. 4(j). Although the occupation probability at the initially localized node decays over time, the distribution remains non-uniform and approaches uniformity much more slowly than in the Erdős-Rényi and small-world networks. The scale-free network shows enhanced robustness under Haken-Strobl noise, preserving coherence for longer times than the other topologies (Fig. 5(c)). In contrast, under QSW decoherence it exhibits the opposite trend. All topologies delocalize and lose coherence significantly faster under QSW decoherence than under the other noise models, as evident from Figs. 4 and 5.
Decoherence Network Single Realization Ensemble
$\lambda$ $\beta$ $\tau$ $\langle\tau\rangle$
Haken-Strobl Erdős-Rényi 0.10 0.84 10.96 11.46
Small-World 0.17 0.96 5.99 11.56
Scale-Free 0.24 0.37 17.43 38.99
Quantum Stochastic
Erdős-Rényi 0.71 0.45 3.49 2.17
Small-World 1.53 0.90 0.69 3.46
Scale-Free 0.96 0.62 1.50 1.72
@@
Table 2: Coherence decay fitting parameters and relaxation time are for Erdős-Rényi, small-world, and scale-free topologies. Values include for both individual realizations ($\tau$) and ensemble-averaged relaxation time ($\langle\tau\rangle$) calculated for 50 realizations under Haken-Strobl and QSW decoherence models.
Further, the decay coefficients $\lambda$ and $\beta$ are obtained by fitting the $\ell_{1}$-norm of coherence (Fig. 5), from which we compute the relaxation time $\tau$ (Table 2) and the ensemble-averaged relaxation time $\langle\tau\rangle$ over $50$ network realizations. These quantitative measures verify our earlier observations that under Haken-Strobl noise, the scale-free networks preserve coherence for longer times than the QSW, as reflected by its comparatively larger values of $\tau$ and $\langle\tau\rangle$.
Conversely, QSW decoherence induces the most rapid decay across all complex topologies, with the scale-free network proving more vulnerable than Erdős–Rényi as shown by a lower value of $\langle\tau\rangle$. Such relative vulnerability is consistent with the findings for simple topologies discussed in Sec. III.1, and stems from the fundamental nature of the type of decoherence mechanisms. Specifically, since QSW jumps are link-dependent, the high degree of the hub nodes in heterogeneous networks causes a higher decoherence and consequently faster decay of coherence. To systematically bridge the network’s structural features with the observed dynamics, Appendix C explores $\langle\tau\rangle$, and the statistical average of initialized node properties. Our results suggest that different properties of initially localized node contribute to capture the actual dynamics of the quantum system under different decoherence models.
In the presence of intrinsic decoherence, the CTQW exhibits partial localization at the initially localized node in the scale-free network, reaching decoherence-induced steady state Fig. 4(k). In contrast, localization behavior of CTQW is not observed in the Erdős–Rényi and small-world networks, which also resist reaching a steady-state within the observed time-scales as illustrated in Fig. 4(c),(g). Interestingly, the $\ell_{1}$-norm of coherence remains preserved under intrinsic decoherence for all network topologies (Fig. 5), consistent with the behavior observed for simple networks in Sec. III.1. This indicates that, although intrinsic decoherence induces partial localization in the scale-free network, it preserves a finite amount of quantum coherence across all topologies, thereby maintaining stability.
@@
Further we use other standard metrics namely quantum-classical distance, von Neumann entropy, and fidelity to check the stability of quantum system. The quantum-classical distance in Fig. 6(a)-(d) suggest that the scale-free network exhibits greater resistance to classicalization under decoherence compared to the Erdős–Rényi and small world network, since as time progresses, scale-free has a higher quantum classical distance than other topologies underlying all decoherence dynamics. It shows a slower decay of quantum-classical distance in the presence of both QSW and Haken-strobl both of which classicalize the system over long timescales. For intrinsic decoherence in Fig. 6(c), the distance does not decay to zero, indicating that classicalization does not occur and quantum features are consistently preserved for all networks. For the noiseless scenario, across all network topologies, the quantum-classical distance converges to the theoretical value $\lim_{t\to\infty}D_{QC}(t)=1-\frac{1}{N}$ [37], as represented by the dashed line in Fig. 6(a).
@@
The fidelity plots in Fig. 6(e)-(h) show that the scale-free network topology exhibits localization near the initial state since it has high fidelity with the initial state. The higher the fidelity lesser the evolved state has deviated from the initial state. On contrast, the Erdős-Rényi and small-world networks have a little less fidelity across various decoherence.
@@
Von Neumann entropy plots Fig. 7(a)-(c) show that entropy is high for Erdős-Rényi in all decoherence. Entropy is least for scale-free in the case of Haken-Strobl and QSW, showing scale-free is not much deviated from the pure state as compared to Erdős-Rényi. This implies that scale-free is more resilient to all noises and Erdős-Rényi is more prone to mixed states caused by decoherence than other topologies. For the noiseless case, the entropy always remains zero.
Taken together, despite the rapid coherence decay under QSW decoherence, scale-free networks consistently maintain higher quantum-classical distance, higher fidelity, and lower entropy, demonstrating their superior stability relative to other topologies.
IV Conclusion and Discussion